# Equidistribution of polynomially bounded o-minimal curves in homogeneous spaces

### v1

@@ 6. Linearization
Let $\mathcal{H}$ be the class of all closed connected subgroups $H$ of $\text{SL}(n,\mathbb{R})$ such that $H/H\cap\Gamma$ admits an $H$-invariant probability measure and the closed subgroup $H_{u}\leq H$ which is generated by all unipotent one-parameter subgroups of $H$ acts ergodically on $H/H\cap\Gamma$ with respect to the $H$-invariant probability measure.
We note the following results:
  • [Ratner91a, Theorem 1.1]: The collection $\mathcal{H}$ is countable.
For $H\in\mathcal{H}$, define
$N(H,W)=$$\{g\in\text{SL}(n,\mathbb{R}):W\subset gHg^{-1}\},\text{ and}$
@@
In what follows, we denote by $\pi:\text{SL}(n,\mathbb{R})\to X$ the canonical projection. Consider
(6.1) $T_{H}(W)=\pi(N(H,W)\smallsetminus S(H,W))=\pi(N(H,W))\smallsetminus\pi(S(H,W)),$
for the second equality see [Mozes1995OnTS, Lemma 2.4].
@@
[Mozes1995OnTS, Theorem 2.2] Let $\nu$ be a $W$-invariant Borel probability measure on $X$. For $H\in\mathcal{H}$, let $\nu_{H}$ denote the restriction of $\nu$ on $T_{H}(W)$. Then $\nu$ decomposes as:
@@
where $\mathcal{H}^{*}\subset\mathcal{H}$ is a countable set of subgroups, each is a representative of a distinct $\Gamma$-conjugacy class.
Moreover, for $H\in\mathcal{H}$ it holds that $\nu_{H}$ is $W$-invariant, and any $W$-ergodic component of $\nu_{H}$ is the unique $gHg^{-1}$-invariant probability measure on the closed orbit $gH\Gamma/\Gamma\subset X$, for some $g\in N(H,W)\smallsetminus S(H,W)$.
@@
For any $d\in\mathbb{N}$, consider the action of $\text{SL}(n,\mathbb{R})$ on
(6.2) $V_{d}=\bigwedge^{d}\mathfrak{sl}(n,\mathbb{R}),$
induced by the Adjoint representation of $\text{SL}(n,\mathbb{R})$ on its Lie algebra $\mathfrak{sl}(n,\mathbb{R})$; that is, $g\cdot(\bigwedge_{i=1}^{d}x_{i}):=\bigwedge_{i=1}^{d}\text{Ad}_{g}(x_{i})$ and extended linearly. Let $H\in\mathcal{H}$, and let $\mathfrak{h}$ be the Lie algebra of $H$ and $d_{H}=\dim\mathfrak{h}$. Fix a vector $p_{H}\in\bigwedge^{d_{H}}\mathfrak{h}\setminus\{0\}$
@@
    [Dani1993LimitDO, Proof of Proposition 3.2] Let $\mathfrak{w}$ be the Lie algebra of $W$, and consider the variety
    (6.3) $A_{H}:=\{v\in V_{d_{H}}:v\wedge w=0\in V_{d_{H}+1},\forall w\in\mathfrak{w}\}.$
    Then, $g.p_{H}\in A_{H}$ if and only if $g\in N(H,W)$. In other words, let $\eta_{H}:\text{SL}(n,\mathbb{R})\to V_{d_{H}}$ denote the orbit map
    (6.4) $\eta_{H}(g)=g.p_{H},\,\forall g\in\text{SL}(n,\mathbb{R}),$
@@
    [Dani1993LimitDO, Theorem 3.4] $\Gamma.p_{H}$ is discrete in $V_{d_{H}}$.
We present the following variant of [Mozes1995OnTS, Proposition 3.4].
@@
Let $\varphi:[0,\infty)\to\text{SL}(n,\mathbb{R})$ be a continuous, unbounded, non-contracting curve that is definable in a polynomially bounded o-minimal structure. Let $\beta:[0,\infty)\to\text{SL}(n,\mathbb{R})$ be a correcting curve of $\varphi$ and let $H_{\varphi}$ be the hull of $\varphi$ such that $\beta(t)\varphi(t)\subset H_{\varphi}$ for all $t$, see Definition 2.5. Fix $H\in\mathcal{H}$. Let $\emph{\text{K}}_{1}\subset A_{H}$ be a compact set, and let $10<\epsilon<$ 1. Then there exists a closed set $\mathcal{S}\subset\pi(S(H,W))$ such that the following holds.
For a given compact set $K\subset X\smallsetminus\mathcal{S}$, there exists a neighborhood $\Psi$ of $\emph{\text{K}}_{1}$ in $V_{d_{H}}$ and such that for any $x\in X$, at least one of the following is satisfied:
@@
    There exists an $w\in\pi^{-1}(x).p_{H}\cap\Psi$ such that
    (6.5) $H_{\varphi}.w=w.$
@@
    For all sufficiently large $T>0$ (depending on $x$),
    (6.6) |$|\{t\in[0,T]:\beta(t)\varphi(t).x\in K\cap\pi(\eta_{H}^{-1}(\overline{\Psi}))\}|\leq\epsilon T.$
    Here, $\eta_{H}$ is the orbit map (6.4).
First, recall some observations and results that will be used in the proof. We let $N(H)$ be the normalizer of $H$ in $\text{SL}(n,\mathbb{R})$, and denote
$N_{\Gamma}:=N(H)\cap\Gamma.$
We recall that for $\gamma\in N_{\Gamma}$ it holds that
(6.7) $\det(\text{Ad}_{\gamma}|_{\mathfrak{h}})\in\{\pm 1\},$
see [Dani1993LimitDO, Lemma 3.1].
In view of (6.7), either $N_{\Gamma}.p_{H}=\{p_{H}\}$ or $N_{\Gamma}.p_{H}=\{\pm p_{H}\}$. If the latter case is true, we will denote by $\tilde{V}_{H}:=V_{d_{H}}/\{\pm 1\}$, the space $V_{d_{H}}$ modulo the equivalence relation defined by identifying $v\in V_{d_{H}}$ with $-v$. Otherwise, namely if $N_{\Gamma}.p_{H}=\{p_{H}\}$, then we denote $\tilde{V}_{d_{H}}:=V_{d_{H}}$.
@@
(6.8) $S(D)=\{g\in\eta_{H}^{-1}(D):g\gamma\in\eta_{H}^{-1}(D)\text{ for some }\gamma\in\Gamma\smallsetminus N_{\Gamma}\}\subset\text{SL}(n,\mathbb{R}).$
@@
    Suppose that $K\subset X\smallsetminus\pi(S(D))$ is compact. Then, there is a compact neighborhood $\Phi\subseteq V_{d_{H}}$ of $D$ such that for every $y\in\pi(\eta_{H}^{-1}({\Phi}))\cap K$, the set $(\pi^{-1}(y).\tilde{p}_{H})\cap\tilde{{\Phi}}$ consists of a single element. Here $\tilde{{\Phi}}$ are the images of $p_{H}$ and ${\Phi}$ in $\tilde{V}_{d_{H}}$.
@@
We assume for simplicity that $N_{\Gamma}.p_{H}=\{p_{H}\}$, namely $\tilde{V}_{d_{H}}=V_{d_{H}}$. The case for which $N_{\Gamma}.p_{H}=\{\pm p_{H}\}$ is treated in essentially the same way and is left for the readers. Let $\varphi:[0,\infty)\to\text{SL}(n,\mathbb{R})$ be an unbounded, continuous non-contracting curve definable in a polynomially bounded o-minimal structure. Fix $H\in\mathcal{H}$. Let $\text{K}_{1}\subset A_{H}$ be a compact subset and $10<\epsilon<$ 1 be given. We will apply Proposition 3.10 with $\psi(t)$ being the image of $\beta(t)\varphi(t)$ in the representation ${V}_{d_{H}}$. Let $\text{K}_{2}\subset A_{H}$ be a compact set given by Proposition 3.10 with $\text{K}_{1}\subset\text{K}_{2}$.
@@
In the notations of Proposition 6.5, we take the compact set $D:=\text{K}_{2}$ and let $\mathcal{S}:=\pi(S(\text{K}_{2})).$ Let $K\subset X\smallsetminus\mathcal{S}$ be a compact set given in the hypothesis of the proposition. Let $\Phi\subset{V}_{d_{H}}$ be a compact neighborhood of $\text{K}_{2}$ which satisfies the outcome of Proposition 6.5,(3). As a consequence, for any $t\geq 0$, $w_{1},w_{2}\in\pi^{-1}(x).p_{H}$, and $g\in\text{SL}(n,\mathbb{R})$, if $gx\in K$ and $\beta(t)\varphi(t).w_{i}\in\Phi$ for $i=1,2$, then $w_{1}=w_{2}$; let us call this the injective property of $K$ in $\Phi$.
We now choose an open neighborhood $\Psi$ of $\text{K}_{1}$ in $V_{d_{H}}$, with $\overline{\Psi}\subset\mathring{\Phi}$, as in Proposition 3.10, where $\mathring{\Phi}$ denotes the interior of $\Phi$.
Case 1There exists $w\in\pi^{-1}(x).p_{H}\cap\overline{\Psi}$ such that $H_{\varphi}.w=w$ for all $t\geq 0$.
Case 2For all $w\in\pi^{-1}(x).p_{H}\cap\overline{\Psi}$ it holds that $\lim_{t\to\infty}\beta(t)\varphi(t).w=\infty$. By Theorem 2.4,(3), the two mutually exclusive cases exhaust all possibilities. To complete the proof, we now assume that Case 1 does not occur and Case 2 occurs. We want to show that (6.6) holds for all sufficiently large $T$.
We take $T^{\prime}_{0}>0$ to satisfy the outcomes of Propositions 1.15 and 3.10, and consider
(6.9) $J=\{t\geq T^{\prime}_{0}:\beta(t)\varphi(t)x\in\pi(\eta^{-1}_{H}(\overline{\Psi}))\cap K\}.$
Then, to prove (6.6), we only need to show that
(6.10) $\lvert J\cap[T_{0}^{\prime},T]\rvert\leq\epsilon T\text{, for all sufficientlylarge $T\geq T_{0}^{\prime}$.}$
For every $t\in J$, there exists $w\in\pi^{-1}(x).p_{H}$ such that $\beta(t)\varphi(t).w\in\overline{\Psi}\subset\Phi$ and we have $\beta(t)\varphi(t)x\in K$. By the injective property of $K$ in $\Phi$, we get that such a $w$ is unique, and denote it by $w(t)$. We note that for any $t\in J$, $\lim_{s\to\infty}\beta(s)\varphi(s).w(t)=\infty$, because otherwise $w(t)$ is fixed by $H_{\varphi}$ due to (3) of Theorem 2.4, and hence $w(t)\in\overline{\Psi}\cap\pi^{-1}(x).p_{H}$, contradicting our assumption that Case 1 does not occur.
Therefore, given $t\in J$ and its corresponding $w(t)\in\pi^{-1}(x).p_{H}$, there exists the largest closed interval $I(w(t)):=[t^{-},t^{+}]\subset[T^{\prime}_{0},T]$ containing $t$ such that
  1. (1)
    $\beta(s)\varphi(s).w(t)\in\Phi$ for all $s\in I(w(t))$,
  2. (2)
    $t^{-}\in J$, and
  3. (3)
    $\beta(t^{+})\varphi(t^{+}).w(t)\in\Phi\smallsetminus\mathring{\Phi}$ or $t^{+}=T$.
We note that $I(w(t))$ is uniquely determined by the $t\in J$. For any $s\in J\cap I(w(t))$, we have $\beta(s)\varphi(s)x\in K$ and $\beta(s)\varphi(s)\cdot w(t)\in\Phi$, so by the injectivity of $K$ in $\Phi$, we have $w(s)=w(t)$. Therefore, by (6.9),
$J\cap I(w(t))$ =$=\{s\in I(w(t))\cap J:\beta(s)\varphi(s)\cdot w(s)\in\Psi\}$
=$=\{s\in I(w(t))\cap J:\beta(s)\varphi(s)\cdot w(t)\in\Psi\}$
@@
(6.11) $\subseteq\{s\in I(w(t)):\beta(s)\varphi(s)\cdot w(t)\in\Psi\}.$
We claim that for any $t_{1},t_{2}\in J$ and their corresponding $w(t_{1}),w(t_{2})\in\pi^{-1}(x).p_{H}$, we have
(6.12) either $\text{either }I(w(t_{1}))=I(w(t_{2}))\text{ or }I(w(t_{1}))\cap I(w(t_{2}))=\varnothing.$
To verify this claim, suppose that $I(w(t_{1}))\cap I(w(t_{2}))\neq\varnothing$. Then either $t_{1}^{-}\in I(w(t_{2}))$ or $t_{2}^{-}\in I(w(t_{1}))$. Since $t_{1}^{-},t_{2}^{-}\in J$, we have $J\cap I(w(t_{1}))\cap I(w(t_{2}))\neq\varnothing$. So, by our earlier observation, $w(t_{1})=w(t_{2})$. This proves the claim.
To prove (6.10), we only need to consider the case of $J$ being unbounded, and $J\cap[T_{0}^{\prime},T]\neq\varnothing$. Let $t_{1}^{-}=\min(J\cap[T_{0}^{\prime},T])$ and $w_{1}=w(t_{1}^{-})=[t_{1}^{-},t_{1}^{+}]$. We inductively define $w_{1},\ldots,w_{l(T)}$ as follows: For any $i\geq 1$ such that $w_{i}$ is defined, if $J\cap(t_{i}^{+},T]\neq\varnothing$, let $t_{i+1}^{-}=\inf(J\cap(t_{i}^{+},T])$, and $w_{i+1}:=w(t_{i+1}^{-})$, and $I(w_{i+1})=[t_{i+1}^{-},t_{i+1}^{+}]$. Otherwise, we let $l(T)=i$ and stop. Thus,
(6.13) $J\cap[T_{0}^{\prime},T]\subseteq I(w_{1})\sqcup...\sqcup I(w_{l(T)}).$
Here among $\{w_{1},\ldots,w_{l(T)}\}\subset\pi^{-1}(x).p_{H}$, each $w_{i}$ appears at most finitely many times, because $\beta(s)\varphi(s)w_{i}\to\infty$ as $s\to\infty$. Since $\pi^{-1}(x).p_{H}$ is a discrete subset of $V_{d_{H}}$, and since and $J$ is unbounded, $\lVert w_{l(T)}\rVert\to\infty$ as $T\to\infty$.
So, by (6.11) and (6.13),
(6.14) $\lvert J\cap[T_{0}^{\prime},T]\rvert\leq\sum_{k=1}^{l(T)}\left|\{t\in I(w_{k}):\beta(t)\varphi(t).w_{k}\in\Psi\}\right|.$
Let $11\leq i\leq l(T)-$ 1. Then $t_{i}^{+}<t_{l_{T}}^{-}<T$. So, by the defining property (3) of $I(w_{i})=[t_{i}^{-},t_{i}^{+}]$, we have $\beta(t_{i}^{+})\varphi(t_{i}^{+})\cdot w_{i}\in\Phi\setminus\mathring{\Phi}$. Now, by Proposition 3.10 we have that
|$\left|\{t\in I(w_{k}):\beta(t)\varphi(t).w_{k}\in\Psi\}\right|\leq\epsilon|I(w_{k})|.$
Therefore,
(6.15) $\sum_{k=1}^{l(T)-1}\left|\{t\in I(w_{k}):\beta(t)\varphi(t).w_{k}\in\Psi\}\right|\leq\epsilon(T-T_{0}).$

### v2

@@ 6.1. Description of Ratner’s singular sets
Let $\mathcal{H}$ be the class of all closed connected subgroups $H$ of $\text{SL}(n,\mathbb{R})$ such that $H/H\cap\Gamma$ admits an $H$-invariant probability measure and the closed subgroup $H_{u}\subseteq H$ which is generated by all unipotent one-parameter subgroups of $H$ acts ergodically on $H/H\cap\Gamma$ concerning the $H$-invariant probability measure.
Theorem 6.2 ([Ratner91a, Theorem 1.1]).
The collection $\mathcal{H}$ is countable.
We note that if $\Gamma=\text{SL}(n,\mathbb{Z})$, then $\text{zcl}(H\cap\Gamma)$ is defined over $\mathbb{Q}$ and $H=\text{zcl}(H\cap\Gamma)(\mathbb{R})^{0}$ [Shah1991, (2.3) and (3.2)] for all $H\in\mathcal{H}$, so $\mathcal{H}$ is countable.
Let $H\in\mathcal{H}$. Define
$N(H,W)=$$\{g\in\text{SL}(n,\mathbb{R}):W\subseteq gHg^{-1}\},\text{ and}$
@@
Note that $N(W)N(H,W)N(H)=N(H,W)$, where $N(F)$ denotes the normalizer of $F$ in $\text{SL}(n,\mathbb{R})$. Also, for all $\gamma\in\Gamma$, we have $\gamma H\gamma^{-1}\in\mathcal{H}$,
$N(H,W)\gamma=N(\gamma^{-1}H\gamma,W)\text{ and }S(H,W)\gamma=S(\gamma^{-1}H\gamma,W).$
Let $\pi:\text{SL}(n,\mathbb{R})\to X=\text{SL}(n,\mathbb{R})/\Gamma$ denote the quotient map. Consider the refined singular set
(6.1) $T_{H}(W)=\pi(N(H,W)\setminus S(H,W))=\pi(N(H,W))\setminus\pi(S(H,W)).$
for the second equality see [Mozes1995OnTS, Lemma 2.4]. Also, $T_{H}(W)=T_{\gamma H\gamma^{-1}}(W)$ for all $\gamma\in\Gamma$.
@@
(Ratner [Ratner91a]) [Mozes1995OnTS, Theorem 2.2] Let $\nu$ be a $W$-invariant Borel probability measure on $X$. For each $H\in\mathcal{H}$, let $\nu_{H}$ denote the restriction of $\nu$ on $T_{H}(W)$, which is $W$-invariant. Then $\nu$ decomposes as:
@@
where $\mathcal{H}^{*}\subseteq\mathcal{H}$ is a countable set of subgroups, each of which is a representative of a distinct $\Gamma$-conjugacy class.
@@ 6.2. O-minimal curves in representations – a dichotomy theorem
For $H\in\mathcal{H}$, consider
(6.2) $V_{H}:=\bigwedge^{\dim(H)}\mathfrak{sl}(n,\mathbb{R})\text{ and }p_{H}\in\bigwedge^{\dim(H)}\mathfrak{h}\setminus\{0\},$
where $\mathfrak{h}$ denotes the Lie algebra of $H$. Consider the action of $\text{SL}(n,\mathbb{R})$ on $V_{H}$ induced by its Adjoint representation its Lie algebra; that is, $g\cdot(\bigwedge_{i=1}^{d}x_{i}):=\bigwedge_{i=1}^{d}\text{Ad}_{g}(x_{i})$ and extended linearly. We denote by $\eta_{H}:\text{SL}(n,\mathbb{R})\to V_{H}$ the orbit map
(6.3) $\eta_{H}(g)=g\cdot p_{H},\,\forall g\in\text{SL}(n,\mathbb{R}).$
The notation $\eta_{H}$ is used only for denoting preimages under the orbit map. We note some basic observations.
@@
    [Dani1993LimitDO, Proof of Proposition 3.2] Let $\mathfrak{w}$ be the Lie algebra of $W$, and consider the linear subspace
    (6.4) $A_{H}:=\{v\in V_{H}:v\wedge w=0,\forall w\in\mathfrak{w}\}.$
@@
    [Dani1993LimitDO, Theorem 3.4]
    $\Gamma\cdot p_{H}$ is closed, and hence discrete, in $V_{H}$.
For example, if $\Gamma=\text{SL}(n,\mathbb{Z})$, then $\mathfrak{h}$ is defined over $\mathbb{Q}$, so we can pick $p_{H}\in\wedge^{d}\mathfrak{h}(\mathbb{Z})\setminus\{0\}$, and hence, $\Gamma\cdot p_{H}\subseteq V_{H}(\mathbb{Z})$.
Next, we will provide a version of [Mozes1995OnTS, Proposition 3.4] for definable curves. Our proof is based on the linearization technique developed by Dani and Margulis [Dani1993LimitDO, Proof of Theorem 1]. The result can be considered an analogue of Theorem 4.2. In what follows, $\mathscr{S}_{\mathrm{pbd}}$ denotes a polynomially bounded o-minimal structure.
@@
Let $\varphi:[0,\infty)\to\text{SL}(n,\mathbb{R})$ be a $\mathscr{S}_{\mathrm{pbd}}$-definable, continuous, unbounded, non-contracting curve. Let $\beta:[0,\infty)\to\text{SL}(n,\mathbb{R})$ be a correcting curve of $\varphi$ and let $H_{\varphi}$ be the hull of $\varphi$ such that $\beta(t)\varphi(t)\subseteq H_{\varphi}$ for all $t$, see Definition 2.6. Fix $H\in\mathcal{H}$. Let $\emph{\text{K}}_{1}\subseteq A_{H}$ be a compact set. Then, given $x\in X$, one of the following holds:
@@
    There exists $w\in\pi^{-1}(x)\cdot p_{H}\cap\emph{\text{K}}_{1}$ such that
    (6.5) $H_{\varphi}\cdot w=\{w\}.$
@@
    Given $10<\epsilon<$ 1, there exists a closed set $\mathcal{S}\subseteq\pi(S(H,W))$ such that given a compact set $K\subseteq X\setminus\mathcal{S}$, there exist a compact neighborhood $\Psi$ of $\emph{\text{K}}_{1}$ in $V_{H}$ and a $T_{x}>0$ such that for all $T>T_{x}$,
    (6.6) |$|\{t\in[0,T]:\beta(t)\varphi(t)x\in K\cap\pi(\eta_{H}^{-1}(\Psi))\}|\leq\epsilon T.$
We now recall some observations and results that will be used in the proof. Let $N(H)$ be the normalizer of $H$ in $\text{SL}(n,\mathbb{R})$. Note that,
(6.7) $g\cdot p_{H}=\det(\text{Ad}_{g}|_{\mathfrak{h}})p_{H},\,\forall g\in N(H).$
Set $N_{\Gamma}:=N(H)\cap\Gamma.$ By [Dani1993LimitDO, Lemma 3.1], $N_{\Gamma}\subseteq\{\pm p_{H}\}$. If $N_{\Gamma}=\{\pm p_{H}\}$, let $\tilde{V}_{H}:=V_{H}/\{\pm 1\}$, the space $V_{H}$ modulo the equivalence relation defined by identifying $v\in V_{H}$ with $-v$. If $N_{\Gamma}\cdot p_{H}=\{p_{H}\}$, let $\tilde{V}_{H}:=V_{H}$.
For any $g\in\text{SL}(n,\mathbb{R})$ and $y=\pi(g)\in X$, we treat $\pi^{-1}(y)\cdot\tilde{p}_{H}=g\Gamma\cdot\tilde{p}_{H}$ as the discrete set of representatives of $y$ in $\tilde{V}_{H}$, where $\tilde{p}_{H}$ denotes the image of $p_{H}$ in $\tilde{V}_{H}$. We need the following result on the uniqueness of representatives in thin neighborhoods.
@@
$S(D)=\{g\in\eta_{H}^{-1}(D):g\gamma\in\eta_{H}^{-1}(D)\text{ for some }\gamma\in\Gamma\setminus N_{\Gamma}\}\subseteq\text{SL}(n,\mathbb{R}).$
@@
    Suppose that $K\subseteq X\setminus\pi(S(D))$ is compact. Then, there is a compact neighborhood $\Phi$ of $D$ in $V_{H}$ such that any $y\in K$ has at most one representative in $\tilde{\Phi}$, where $\tilde{{\Phi}}$ is the image of ${\Phi}$ in $\tilde{V}_{H}$.
    That is, for all $y\in\pi(\eta_{H}^{-1}({\Phi}))\cap K$, the set $\pi^{-1}(y).\tilde{p}_{H}\cap\tilde{{\Phi}}$ consists of a single element.
@@
We assume for simplicity that $N_{\Gamma}\cdot p_{H}=\{p_{H}\}$, namely $\tilde{V}_{H}=V_{H}$. The case for which $N_{\Gamma}\cdot p_{H}=\{\pm p_{H}\}$ can be treated in essentially the same way, and is left to the readers.
By choosing a basis for $V_{H}$, we identify $V_{H}\cong\mathbb{R}^{m}$ for some $m$. We will denote by $\psi(t)\in\text{GL}(m,\mathbb{R})$ the action of $\beta(t)\varphi(t)$ on $V_{H}$. We will apply Proposition 3.12 for $\psi(t)$. Since $A_{H}$ is a linear subspace, $A_{H}=\{Q=0\}$ for a polynomial map $Q:V_{H}\to\mathbb{R}$, see Remark 3.14. Let $K_{1}\subseteq A_{H}$ and let $10<\epsilon<$ 1 be given. Let $\text{K}_{2}\subseteq A_{H}$ be a compact set that is given by Proposition 3.12 with $\text{K}_{1}\subseteq\text{K}_{2}$.
Choices of $\mathcal{S}$ and $\Phi$ for injective property
@@
For applying Proposition 6.6, we use $\text{K}_{2}$ in place of $D$, and let $\mathcal{S}:=\pi(S(\text{K}_{2}))$. Let $K\subseteq X\setminus\mathcal{S}$ be a compact set, and let $\Phi\subseteq V_{H}$ be a compact neighborhood of $\text{K}_{2}$ which satisfies the outcome of Proposition 6.6,(3). As a consequence, for any $t\geq 0$, $w_{1},w_{2}\in\Gamma\cdot p_{H}$, and $g\in\text{SL}(n,\mathbb{R})$, if $g\Gamma\in K$ and $g\cdot w_{i}\in\Phi$ for $i=1,2$, then $w_{1}=w_{2}$; let us call this the injective property of $K$ in $\Phi$.
Choice of $\Psi$ using $(C,\alpha)$-good property
We take $T^{\prime}_{0}>0$ to satisfy the outcomes of Propositions 1.13 and 3.12 for $\psi$. By Proposition 3.12, we choose a compact neighborhood $\Psi$ of $\text{K}_{1}$ in $V_{H}$ such that $\Psi\subseteq\mathring{\Phi}$ and for any $v\in V$ and any closed interval $[a,b]\subseteq[T_{0}^{\prime},\infty)$, if $\psi([a,b])v\subseteq\Phi$ and $\psi(b)v\notin\mathring{\Phi}$, then
(6.8) |$\lvert\{s\in[a,b]:\psi(s)v\in\Psi\}\rvert\leq\epsilon(b-a).$
It may be noted that in view of (3.30) and (3.31), there exist $R_{2}>0$ and $\epsilon_{1}>0$ such that
(6.9) $\Psi\subseteq\{v\in V_{H}:\lVert v\rVert<R_{2}\text{ and }\lvert Q(v)\rvert<\epsilon_{1}\}\subseteq\mathring{\Phi}.$
Modifying $\Psi$ for dichotomy
Let $x\in X$ be given. Since $\pi^{-1}(x)\cdot p_{H}\cap\Psi$ is finite, by replacing $\Psi$ with a smaller compact neighborhood of $\text{K}_{1}$, we may assume that $\pi^{-1}(x)\cdot p_{H}\cap\Psi=\pi^{-1}(x)\cdot p_{H}\cap K_{1}$.
By Theorem 2.5,(3), the following two mutually exclusive cases exhaust all possibilities.
  • Case 1
    There exists $w\in\pi^{-1}(x)\cdot p_{H}\cap\Psi$ such that $H_{\varphi}\cdot w=\{w\}$.
  • Case 2
    For all $w\in\pi^{-1}(x)\cdot p_{H}\cap\Psi$, $\lim_{s\to\infty}\psi(s)w=\infty$.
If Case 1 holds, then by our choice of $\Psi$, the first conclusion of the theorem holds. Therefore, to complete the proof, we assume that Case 2 holds.
Consider
(6.10) $J=\{t\geq T^{\prime}_{0}:\beta(t)\varphi(t)x\in\pi(\eta^{-1}_{H}(\Psi))\cap K\}.$
Then $J$ is closed. Because $\pi(\eta^{-1}_{H}(\Psi))$ is closed in $X$. To show this, suppose $g_{i}\to g$ in $\text{SL}(n,\mathbb{R})$ and $\gamma_{i}\in\Gamma$ are such that $g_{i}\gamma_{i}\cdot p_{H}\in\Psi$. Since $\Psi$ is compact and $\Gamma\cdot p_{H}$ is discrete, we can pick $\gamma\in\Gamma$ such that $\gamma_{i}\cdot p_{H}=\gamma\cdot p_{H}$ for infinitely many $i$’s. Therefore $g\gamma\cdot p_{H}\in\Psi$.
Then, to prove (6.6), it is enough to show that there exists $T_{x}\geq T_{0}^{\prime}$ depending on $x$ such that
(6.11) $\lvert J\cap[T_{0}^{\prime},T]\rvert\leq 2\epsilon T,\ \forall\,T\geq T_{x}.$
Without loss of generality, we may assume that $J$ is unbounded.
Choice of a representative of $x$ in for every $t\in J$
For every $t\in J$, there exists a representative of $x$, say $w\in\pi^{-1}(x)\cdot p_{H}$, such that $\psi(t)w\in\Psi\subseteq\Phi$. Since $\beta(t)\varphi(t)x\in K$, by the injective property of $K$ in $\Phi$, $w$ is unique, and we denote it by $w_{t}$.
Let $t\in J$. We note that
$\lim_{s\to\infty}\psi(s)w_{t}=\infty,$
because otherwise $w_{t}$ is fixed by $H_{\varphi}$ due to (3) of Theorem 2.5, and since $\psi(t)\subseteq H_{\phi}$, we have $w_{t}\in\Psi\cap\pi^{-1}(x)\cdot p_{H}$, contradicting the Case 2.
Therefore, $\{s\geq T_{0}^{\prime}:\psi(s)w_{t}\in\mathring{\Phi}\}$ is a bounded open subset of $[T_{0}^{\prime},\infty)$ containing $t$. Let $U(t)$ be its connected component containing $t$, and denote
$t^{+}:=\sup U(t)<\infty.$
Then
(6.12) $\psi([t,t^{+}))w_{t}\subseteq\mathring{\Phi}\text{ and }\psi(t^{+})w_{t}\in\Phi\setminus\mathring{\Phi}.$
Bound on $|J\cap[t,t^{+}]|$
@@
For any $s\in J\cap[t,t^{+}]$, we have $\beta(s)\varphi(s)x\in K$, $\beta(s)\varphi(s)\cdot w_{s}\in\Psi\subseteq\mathring{\Phi}$, and $\beta(s)\varphi(s)\cdot w_{t}\in\Phi$, and hence by the injectivity of $K$ in $\Phi$, $w_{s}=w_{t}$. Then,
(6.13) $\begin{array}[]{ll}J\cap[t,t^{+}]&=\{s\in[t,t^{+}]\cap J:\psi(s)w_{s}\in\Psi\}\\ &=\{s\in[t,t^{+}]\cap J:\psi(s)w_{t}\in\Psi\}.\end{array}$
By (6.8), (6.12), and (6.13),
(6.14) $\lvert J\cap[t,t^{+}]\rvert\leq\left|\{s\in[t,t^{+}]:\psi(s)w_{t}\in\Psi\}\right|\leq\epsilon(t^{+}-t).$
Choosing a sequence $t_{i}\in J$
For each $i\in\mathbb{N}$, we define $t_{i}\in J$ inductively as follows: Let $t_{1}=\inf J\in J$. Let $i\in\mathbb{N}$ be such that $t_{i}$ is defined. Then we let $t_{i+1}=\inf J\cap(t_{i}^{+},\infty)$. Since $J$ is closed, $t_{i}\in J$.
By our construction, $t_{i}<t_{i}^{+}<t_{i+1}$ for all $i\in\mathbb{N}$. Let $i\in\mathbb{N}$. Since
$J\cap[T_{0}^{\prime},t_{i}^{+}]=\cup_{j=1}^{i}J\cap[t_{j},t_{j}^{+}),$
by (6.14),
(6.15) $\lvert J\cap[T_{0}^{\prime},t_{i}^{+}]\rvert\leq\epsilon(t_{i}^{+}-T_{0}^{\prime}).$
We will bound $|J\cap[t_{i+1},T]|$ using Proposition 1.13. We need the following:
Claim.
As $i\to\infty$, $\lVert w_{t_{i}}\rVert\to\infty$ and $t_{i}\to\infty$.
For $w\in\pi^{-1}(x)\cdot p_{H}$, the maps $s\mapsto\lVert\psi(s)w\rVert$ and $s\mapsto Q(\psi(s)w)$ are definable and for $s\geq T_{0}^{\prime}$. Let
$S_{w}:=\{s\geq T_{0}^{\prime}:\lVert\psi(s)w\rVert=R_{2}\text{ or }\lvert Q(\psi(s)w)\rvert=\epsilon_{1}\}.$
Since $S_{w}$ is a definable subset of $\mathbb{R}$, it has only finitely many connected components, see Theorem B.6. For any $i\in\mathbb{N}$, if $w=w_{t_{i}}$, then $(t_{i},t_{i}^{+})$ contains at least one connected component of $S_{w}$; indeed, since $\psi(t_{i})w_{t_{i}}\in\Psi$ and $\psi(t_{i}^{+})w_{t_{i}}\not\in\mathring{\Phi}$, by (6.9), $t_{i},t_{i}^{+}\not\in S_{w}$ and $(t_{i},t_{i}^{+})\cap S_{w}\neq\emptyset$. Therefore, since $\{t_{i}\}_{i=1}^{\infty}$ is an infinite sequence, the set $\{i\in\mathbb{N}:w=w_{t_{i}}\}$ is finite.
So, $\{w_{t_{i}}:i\in\mathbb{N}\}$ is infinite. Since $\pi^{-1}(x)\cdot p_{H}$ is discrete in $V_{H}$, we get $\lim_{i\to\infty}\|w_{t_{i}}\|=\infty$. Since $\psi(t_{i})w_{t_{i}}\in\Psi$ for all $i$ and the map $s\mapsto\lVert\psi(s)^{-1}\rVert$ is continuous on $[T_{0}^{\prime},\infty)$, we conclude that $t_{i}\to\infty$. This proves the claim.